# Are Generative AI systems Capable of Supporting Information Needs of Patients?

### v1

@@ Abstract.
Patients managing a complex illness such as cancer face a complex information challenge where they not only must learn about their illness but also how to manage it. Close interaction with healthcare experts (radiologists, oncologists) can improve patient learning and thereby, their disease outcome. However, this approach is resource intensive and takes expert time away from other critical tasks. Given the recent advancements in Generative AI models aimed at improving the healthcare system, our work investigates whether and how generative visual question answering systems can responsibly support patient information needs in the context of radiology imaging data. We conducted a formative need-finding study in which participants discussed chest computed tomography (CT) scans and associated radiology reports of a fictitious close relative with a cardiothoracic radiologist. Using thematic analysis of the conversation between participants and medical experts, we identified commonly occurring themes across interactions, including clarifying medical terminology, locating the problems mentioned in the report in the scanned image, understanding disease prognosis, discussing the next diagnostic steps, and comparing treatment options. Based on these themes, we evaluated two state-of-the-art generative visual language models against the radiologist’s responses. Our results reveal variability in the quality of responses generated by the models across various themes. We highlight the importance of patient-facing generative AI systems to accommodate a diverse range of conversational themes, catering to the real-world informational needs of patients.
Generative AI, Healthcare, Medicine, LLM, Visual Question Answering
copyright: noneccs: Human-centered computing User studiesccs: Computing methodologies Knowledge representation and reasoningccs: Computing methodologies Natural language processing
1. Introduction
@@
Patients who are better educated about their disease have better outcomes, especially in serious illnesses such as cancer  (adams2010improving, ; berkman2011, ; howell2017self, ; christiansen2023systematic, ). However, they and their caregivers face complex information and learning challenges when navigating healthcare systems. They not only must learn what clinical terminology means but also what their medical reports convey, how the severity of their disease is diagnosed, what their treatment options are etc. This often leads to anxiety and stress from not only from the disease but also from the complexity of understanding and managing it. To alleviate these challenges, healthcare systems have invested significantly in patient education resources. A crucial resource is meaningful and empathetic interactions with healthcare providers supporting the patient’s journey through the healthcare system (gilligan2018patient, ). While arguably the most useful, it also takes the experts’ time away from their other critical responsibilities, adding to their time burden. On the other hand, patients often desire additional informational support and augment their understanding with generic and unverified information from search engines (de2019quality, ) and forums (alarifi2021understanding, ).
Recently, conversational generative AI systems built upon large language models and their multi-modal counterparts have shown great promise as tools that enable people to access information through natural conversations (dwivedi2023so, ). Generative models for medicine such as MedFlamingo (moor2023med, ), GPT-4v (yang2023dawn, ), LLava-Med (li2023llava, ) have shown impressive performance in visual question answering tasks comprising of medical scans including Computed Tomograhpy Scan (CT scan) and accompanying medical reports  (wu2023towards, ). Purportedly, these models have medical common sense at par with doctors as evidenced by their high scores in standard medical knowledge evaluation paradigms like the United States Medical Licensing Examination (USMLE) (jin2021disease, ). However, unlike usecases involving factual knowledge for experts, integrating generative models in patient-facing tools requires that the models also consider teaching and explanation of diagnosis, and prognosis; weigh risks, benefits, and costs of treatments; and understand complex human emotional and social contexts. Lack of adequate considerations can lead to harmful consequences for patients, including misinterpretation of generated output, over-reliance on AI, and poor disease management. Our goal is assess whether generative AI systems can support patients in their informational needs as they learn more about their illness and reason about next steps.
@@
In this work, we study how caregivers interact with radiologists to gain a better understanding of their scans and reports. This use case is relevant for multi-modal generative AI systems that can combine the visual channel of scan images and the linguistic channel of reports. In collaboration with a cardiothoracic radiologist at a major university hospital (and a co-author), we conducted a formative need-finding study with $5$ participants. The study was framed as a conversational interview in which participants discussed chest computed tomography (CT) scans and associated radiology reports of a fictitious close relative with a radiologist. We limited the participants to a small number because the subject of the study is highly sensitive and could trigger anxiety in our participants. Additionally, each interview required ${}3hrs$ of the radiologist’s time. We analyzed the recorded interviews with inductive thematic analysis to identify emerging themes and categories in the conversations. Next, we evaluated two multi-modal generative AI systems - MedFlamingo (moor2023med, ) and ChatGPT-4V (yang2023dawn, ) - against the radiologist’s responses in the realistic interview task across four high-level themes of conversation identified. Our findings highlight the strengths and weaknesses of generative AI models in addressing the information needs of patients and caregivers.
@@
Our work situates and evaluates generative AI capabilities in a specific human context of seeking information to enhance understanding in healthcare settings. This enables us to identify the variety of information patients seek and study if AI systems are up to the challenge. Further, generative AI systems are typically evaluated on datasets curated by and for clinicians. The datasets contain questions and answers that are most interesting to clinicians, while performance measurement includes criteria such as diagnostic accuracy. Such an approach introduces an implicit bias in the design and analysis of these systems - they are designed to replace clinicians by mimicking the output of their problem-solving process, following the AI-is-automation philosophy. Instead, our research takes the AI-is-augmentation view and centers patients and caregivers, adopting a much-needed lens missing from the medical generative AI research landscape. We introduce a novel qualitative answer evaluation paradigm that measures both factual correctness and relevance of answers generated by AI systems. Our research lays the groundwork for a rigorous, quantitative approach to measuring the suitability of a generative AI system for addressing patients’ informational needs. Our key contributions include: (1) a set of conversational themes central to patient-facing LLM systems and (2) an evaluation of quality of responses from two state-of-the-art medical generative models.
2. Background
@@
Recently, large language models (LLMs) like PaLM (chowdhery2023palm, ), GPTs (radford2019language, ; brown2020language, ; openai2023gpt4, ), LLaMA (touvron2023llama, ; touvron2023llama2, ) have significantly advanced the state-of-the-art in various natural language processing (NLP) tasks (zhao2023survey, ; yang2023harnessing, ). This has inspired the development of many medical LLMs (singhal2023large, ; singhal2023towards, ; han2023medalpaca, ; toma2023clinical, ; yunxiang2023chatdoctor, ; liu2023medical, ) to assist the healthcare professionals and improve patient care (arora2023promise, ; thirunavukarasu2023large, ; patel2023chatgpt, ). Based on publicly available LLMs, specialized medical versions of the models like ChatDoctor (yunxiang2023chatdoctor, ), MedAlpacha (han2023medalpaca, ), PMC-LLaMA (wu2023pmc, ), BenTsao (wang2023huatuo, ), and Clinical Camel (toma2023clinical, ) have come up in the past year. Models like Med-PaLM (singhal2023large, ) and Med-PaLM-2 (singhal2023towards, ) built on top of PaLM have achieved close to human expert scores in the United States Medical Licensing Examination (USMLE) (jin2021disease, ). Apart for only text-based models several multimodal models (liu2023visual, ; awadalla2023openflamingo, ; li2023blip, ; yang2023dawn, ) have come to prominence in the last year with the capability of processing both images, text and videos to generate responses. Similar to NLP models such multimodal models have promising application in the medical domain. They can be used as visual question answering systems for patient medical reports, medical report summarization and analysis tasks among many other possibilities. In this work, we utilize MedFlamingo (moor2023med, ) and GPT-4V (yang2023dawn, ) as our generative AI models to evaluate human expert responses in patient-radiologist interactions.
@@
MedFlamingo (moor2023med, ) leverages the Flamingo architecture (alayrac2022flamingo, ) as its foundation. It is pre-trained on paired medical image-text data from medical publications and textbooks. This multimodal few-shot learner has been demonstrated to have remarkable performance in medical visual question answering (VQA) tasks, surpassing existing models by up to 20% in clinician evaluation score as reported in the original work. (moor2023med, ) Unlike previous efforts in multimodal medical foundation models, such as ChexZero(tiu2022expert, ) and BiomedCLIP(zhang2023large, ), Med-Flamingo is designed specifically for in-context learning in the medical domain and can learn tasks from few examples during prompting. More recently, GPT-4 model with vision capabilities (GPT-4V) (yang2023dawn, ) has been identified as the most advanced LLM thus far, particularly in its application to radiology domain, where it has been compared against the current state-of-the-art (SOTA) models(liu2023exploring, ). The research, which referenced various sources, thoroughly examined GPT-4V across a broad spectrum of standard radiology text-based tasks, including MS-CXR-T(bannur2023mscxr, ) and CheXbert(smit2020chexbert, ). The findings revealed that GPT-4 either surpasses or matches the performance of specialized fine-tuned radiology models. A detailed analysis in a Microsoft explorations paper(yang2023dawn, ) on GPT-4V specifically investigated its utility in generating radiology reports. This investigation demonstrated that GPT-4V could accurately diagnose and recommend management based on X-ray images. The findings from this research indicate that GPT-4V holds significant promise as an AI assistant in radiology report generation.
@@
Despite the growing interest in the application of Generative AI models in the medical and healthcare domains, there are a few key issues that need to be addressed for their reliable and effective usage in such a critical domain. Several recent surveys (zhao2023survey, ; he2023survey, ; xie2023faithful, ) on the application of LLMs for medicine highlight the best practices and challenges involved in their application and development. Effective evaluation of AI systems build around LLMs is one such challenge. Particularly in the medical domain, the current benchmarks and metrics often fail to evaluate LLMs’ overall capabilities and emerging abilities. Current benchmarks such as USMLE MedQA (jin2021disease, ) and MedMCQA (pal2022medmcqa, ) cover broad range of question-answering tasks but lack the information to evaluate on metrics like trustworthiness and helpfulness to different parties involved like patients, doctors and administrators (xie2023faithful, ). It is imperative to prepare more domain and task specific evaluation protocols to adjudicate the properties of GenAI models. Singhal et al. (singhal2023large, ) created a benchmark comprising the most commonly searched health queries to evaluate LLM responses. TruthfulQA (lin2021truthfulqa, ) and HaluEval (li2023halueval, ) attempt to evaluate truthfulness and factual accuracy, but not for any specific medical domain. The analyses reported in our work is the first attempt to reflect upon the conversational patterns and evaluate strengths and weakness of generative AI system responses against human experts in a real-world patient-doctor interaction.
3. Preliminaries
@@
The average patient’s journey through the medical system in the United States is shown in Figure 1. Typically, a patient first enters the medical system and meets with a primary care physician (PCP) who collects their history, performs a physical evaluation, and orders further medical and radiology-based tests in order to inform a diagnosis. A radiologic technologist performs the radiology tests the referring PCP orders, and passes on these scans to the consulting radiologist. The radiologist analyses these scans and generates an associated radiology report. The referring physician receives these scans, reports, and the results of other medical tests they had ordered, and combines this information to arrive at an appropriate diagnosis. They then meet with the patient again to discuss the diagnosis, and followup and treatment options. It is relatively rare for radiologists to directly meet with their patients to help them understand their scans and radiology reports (glazer2011invisible, ; kemp2017patient, ).
@@
Figure 1. Patient journey through the medical system highlight various interactions they have with different healthcare experts.
@@
Since the radiology report is intended to aid the patient’s physician in arriving at a diagnosis, and not the patient themselves, the report contains plenty of medical terminology that is not accessible to the average patient. However, the patient receives immediate access to their radiology report once it is generated, in their electronic health records. Since it is atypical and oftentimes difficult for a patient to schedule a one-on-one consultation with their consulting radiologist to understand their scan and the associated report, patients have to wait two to three weeks on average to meet with their referring physician instead. This can tend to create anxiety in patients faced with medical information they cannot understand. Often, patients resort to unstructured, unverified resources such as Reddit for further information.
An effective multi-modal generative AI system can be useful during the period that the patient is waiting to meet with their PCP. The AI system can support a patient understand medical terminology used in the report and how the terms relate to the scan. It can help the patient understand the causes of their disease and its prognosis. It can support patients research various treatment options and weight their risks and benefits. Engaging with an intelligent information tool can prepare the patient (and their caregivers) for a productive meeting with their PCP.
4. Formative Study of Patient-Provider Interaction
@@
4.1. Study
Figure 2. 2 axial slices of the lung CT and the radiology report presented to study participants
4.1.1. Materials
@@
4.1.2. Protocol
Figure 3 shows the study protocol we implemented. The study progressed as follows:
  1. (1)
@@ 4.1.2. Protocol
Figure 3. Protocol showing the context and task prompt as well as participant actions and interactions
4.1.3. Conducting the study
@@
4.2. Inductive Thematic Analysis
We transcribed the interaction sessions using Zoom’s transcription service. Our final dataset consisted of $5$ documents averaging $3049$ ($R=1642-4614$) words. We analyzed this data set with inductive thematic analysis (nowell2017thematic, ). We started with a single interview and went through the transcript one sentence at a time, assigning codes that captured the content of the sentence, or a set of related sentences. For each new interview we analyzed, we used existing codes generated based on previous interviews and assigned new codes only when there were no existing codes to capture the content of a specific statement or set of statements. Each time a new set of codes was added based on an interview, we revisited previously coded interviews to check if the new codes could be assigned to any statements in them. We found that we reached data saturation at $5$ interviews with only $2$ new codes being added in analysing the final interview. We identified a set of $91$ content-based codes across all transcribed participant-radiologist interactions, and we grouped these codes into a set of $10$ broad themes that best captured the relationship between the codes.

### v2

# Joint Proceedings of the ACM IUI Workshops 2025, March 24-27, 2025, Cagliari, Italy

@@ Abstract
Generative AI systems such as ChatGPT and Claude are built upon language models that are typically evaluated for accuracy on curated benchmark datasets. Such evaluation paradigms measure predictive and reasoning capabilities of language models but do not assess if they can provide information that is useful to people. In this paper, we take some initial steps in developing an evaluation paradigm that centers human understanding and decision-making. We study the utility of generative AI systems in supporting people in a concrete task - making sense of clinical reports and imagery in order to make a clinical decision. We conducted a formative need-finding study in which participants discussed chest computed tomography (CT) scans and associated radiology reports of a fictitious close relative with a cardiothoracic radiologist. Using thematic analysis of the conversation between participants and medical experts, we identified commonly occurring themes across interactions, including clarifying medical terminology, locating the problems mentioned in the report in the scanned image, understanding disease prognosis, discussing the next diagnostic steps, and comparing treatment options. Based on these themes, we evaluated two state-of-the-art generative AI systems against the radiologist’s responses. Our results reveal variability in the quality of responses generated by the models across various themes. We highlight the importance of patient-facing generative AI systems to accommodate a diverse range of conversational themes, catering to the real-world informational needs of patients.
keywords:
[…]
1 Introduction
Conversational systems built upon generative AI technology are general-purpose tools that enable people to access information through natural conversations  [1]. Healthcare is primed to have significant benefits with reliable, robust and trustworthy adoption of such systems, assisting different medical interactions/conversations involving patients, doctors, clinicians, and healthcare administrators. Purportedly, generative AI systems have medical commonsense at par with doctors as evidenced by their high scores in standard medical knowledge evaluation paradigms like the United States Medical Licensing Examination (USMLE) [2]. Despite impressive results on benchmark medical and clinical datasets, the case for their adoption in real world clinical scenarios is still unclear. While several clinical researchers and practitioners have expressed excitement about their potential in healthcare [3, 4], several critical questions must be answered before their adoption can begin. These questions include what role generative AI systems can play in a healthcare system, how their behavior should be modulated, and how their performance can be evaluated. Answering these questions requires going beyond benchmark datasets prevalent in AI & ML research to understand the nature of real world problems and characteristics of desired solutions.
@@ 1 Introduction
In this paper, we study a real world clinical usecase for generative AI systems. Patients (along with their caregivers) who are better educated about their disease have better outcomes, especially in serious illnesses such as cancer  [5, 6, 7, 8]. However, they face complex information and learning challenges when navigating healthcare systems. They not only must learn what clinical terminology means but also what their medical reports convey, how the severity of their disease is diagnosed, what their treatment options are etc. This often leads to anxiety and stress not only from the disease but also from the complexity of understanding and managing it. To alleviate these challenges, healthcare systems have invested significantly in patient education resources. Meaningful and empathetic interactions with healthcare providers has a significant, positive impact on a patient’s journey through the healthcare system [9]. While arguably the most useful, it also takes the experts’ time away from their other critical responsibilities, adding to their time burden. On the other hand, patients often desire additional informational support and augment their understanding with generic and unverified information from search engines [10] and forums [11]. Integrating generative AI systems in patient-facing tools requires that generative AI systems be designed to teach and explain diagnosis, and prognosis; weigh risks, benefits, and costs of treatments; and understand complex human emotional and social contexts. Lack of adequate considerations can lead to harmful consequences for patients, including misinterpretation of generated output, over-reliance on AI, and poor disease management.
@@
We examine the scenario where patients and caregivers attempt to understand their medical scans and reports by interacting with a clinical expert. In collaboration with a cardiothoracic radiologist at a major university hospital (and a co-author), we conducted a formative need-finding study with $5$ participants. The study was framed as a conversational interview in which participants discussed chest computed tomography (CT) scans and associated radiology reports of a fictitious close relative with a radiologist. We limited the participants to a small number because the subject of the study is highly sensitive and could trigger anxiety in our participants. Additionally, each interview required ${}3hrs$ of the radiologist’s time. We analyzed the recorded interviews with inductive thematic analysis to identify emerging themes and categories in the conversations. Next, we evaluated two state-of-art multi-modal generative AI systems ChatGPT-4o [12] and Claude 3.5 Sonnet [19] - against the radiologist’s responses in the realistic interview task. Our findings highlight the strengths and weaknesses of generative AI models in addressing the information needs of patients and caregivers.
@@
Our work situates and evaluates generative AI capabilities in a specific human context of seeking information to enhance understanding. This context enables us to identify the variety of information patients seek and study if AI systems are up to the challenge. Further, generative AI systems are typically evaluated on datasets curated by and for clinicians. These datasets contain questions and answers that are most interesting to clinicians, while performance measurement includes criteria such as diagnostic accuracy. Such an approach introduces an implicit bias in the design and analysis of these systems - they are designed to replace clinicians by mimicking the output of their problem-solving process, following the AI-is-automation philosophy. Instead, our research takes the AI-is-augmentation view and centers patients and caregivers. We explore if generative AI systems can support humans - patients and caregivers - in a specific task - making sense of the their reports and scans and accessing knowledge about their disease and treatment.
Advancing the science of human-centered, task-centered evaluation of generative AI systems, our paper makes the following contributions. It,
  • Identifies an opportunity for adoption of generative AI system in clinical practice and specifies an AI task reflecting a real world problem;
  • Characterizes the variability in types of information people seek when making sense of medical reports and scans;
  • Proposes an evaluation approach for conversational systems quantitatively using semantic similarity as well as qualitatively, building upon the Gricean maxims of effective communication;
  • Evaluates two state-of-art generative AI systems - ChatGPT-4o and Claude 3.5 Sonnet - on the proposed evaluation rubric.
2 Background
@@
Recently, large language models (LLMs) like PaLM [13], GPTs [14, 15, 16], LLaMA [17, 18], and Claude [19] have significantly advanced the state-of-the-art in various natural language processing (NLP) tasks [20, 21]. This has inspired the development of many medical LLMs [22, 23, 24, 25, 26, 27] to assist the healthcare professionals and improve patient care [28, 29, 30]. Based on publicly available LLMs, specialized medical versions of the models like ChatDoctor [26], MedAlpacha [24], PMC-LLaMA [31], BenTsao [32], and Clinical Camel [25] have come up in the past years. Models like Med-PaLM [22] and Med-PaLM-2 [23] built on top of PaLM have achieved close to human expert scores in the United States Medical Licensing Examination (USMLE) [2]. Apart for only text-based models several multimodal models [33, 34, 35, 12] have come to prominence in the last year with the capability of processing both images, text and videos to generate responses. Similar to NLP models such multimodal models have promising application in the medical domain. They can be used as visual question answering systems for patient medical reports, medical report summarization and analysis tasks among many other possibilities. In this work, we evaluate Claude [19] and GPT-4V [12] in a patient-facing, clinical usecase.
@@
GPT-4 model with vision capabilities (GPT-4V) [12] has been identified as the most advanced LLM thus far, particularly in its application to radiology domain, where it has been compared against the current state-of-the-art (SOTA) models [36]. The research, which referenced various sources, thoroughly examined GPT-4V across a broad spectrum of standard radiology text-based tasks, including MS-CXR-T[37] and CheXbert[38]. The findings revealed that GPT-4 either surpasses or matches the performance of specialized fine-tuned radiology models. A detailed investigation [12] of GPT-4V studied its utility in generating radiology reports. This investigation demonstrated that GPT-4V could accurately diagnose and recommend management based on X-ray images. The findings from this research indicate that GPT-4V holds significant promise as an AI assistant in radiology report generation. Claude [19] models are more recent but boast multi-modal reasoning capabilities at par to GPT and were the top-performing models on a recent evaluation on medical, multi-modal answer generation benchmarks [39].
@@
Despite the growing interest in the application of Generative AI models in the medical and healthcare domains, there are a few key issues that need to be addressed for their reliable and effective usage in such a critical domain. Several recent surveys [20, 40, 41] on the application of LLMs for medicine highlight the best practices and challenges involved in their application and development. Effective evaluation of AI systems build around LLMs is one such challenge. Particularly in the medical domain, the current benchmarks and metrics often fail to evaluate LLMs’ overall capabilities and emerging abilities. Current benchmarks such as USMLE MedQA [2] and MedMCQA [42] cover broad range of question-answering tasks but lack the information to evaluate on metrics like trustworthiness and helpfulness to different parties involved like patients, doctors and administrators [41]. It is imperative to prepare more domain and task specific evaluation protocols to adjudicate the properties of GenAI models. Singhal et al. [22] created a benchmark comprising the most commonly searched health queries to evaluate LLM responses. TruthfulQA [43] and HaluEval [44] attempt to evaluate truthfulness and factual accuracy, but not for any specific medical domain. The analyses reported in our work is the first attempt to reflect upon the conversational patterns in real-world patient-doctor interactions and evaluate strengths and weakness of generative AI system responses against those of human experts.
3 Preliminaries
@@
The average patient’s journey through the medical system in the United States is shown in Figure 1. Typically, a patient first enters the medical system and meets with a primary care physician (PCP) who collects their history, performs a physical evaluation, and orders further medical and radiology-based tests in order to inform a diagnosis. A radiologic technologist performs the radiology tests the referring PCP orders, and passes on these scans to the consulting radiologist. The radiologist analyses these scans and generates an associated radiology report. The referring physician receives these scans, reports, and the results of other medical tests they had ordered, and combines this information to arrive at an appropriate diagnosis. They then meet with the patient again to discuss the diagnosis, and followup and treatment options. It is relatively rare for radiologists to directly meet with their patients to help them understand their scans and radiology reports [45, 46].
@@
Figure 1: Patient journey through the medical system highlight various interactions they have with different healthcare experts.
@@
Since the radiology report is intended to aid the patient’s physician in arriving at a diagnosis, and not the patient themselves, the report contains medical terminology that is not accessible to the average patient. However, the patient receives immediate access to their radiology report once it is generated, in their electronic health records. Since it is atypical and oftentimes difficult for a patient to schedule a one-on-one consultation with their consulting radiologist to understand their scan and the associated report, patients have to wait two to three weeks on average to meet with their referring physician instead. This can tend to create anxiety in patients faced with medical information they cannot understand. Often, patients resort to unstructured, unverified resources such as Reddit for further information.
An generative AI system can be useful during the period that the patient is waiting to meet with their PCP. The AI system can support a patient understand medical terminology used in the report and how the terms relate to the scan. It can help the patient understand the causes of their disease and its prognosis. It can support patients research various treatment options and weight their risks and benefits. Engaging with an intelligent information tool can prepare the patient (and their caregivers) for a productive meeting with their PCP.
4 Formative Study of Patient-Provider Interaction
@@
4.1 Study
Figure 2: 2 axial slices of the lung CT and the radiology report presented to study participants
4.1.1 Materials
@@
4.1.2 Protocol
Figure 3 shows the study protocol we implemented. The study progressed as follows (item numbers below correspond to stages of the study depicted in 3):
  1. 1.
@@ 4.1.2 Protocol
Figure 3: Protocol showing the context and task prompt as well as participant actions and interactions
4.1.3 Conducting the study
@@
4.2 Inductive Thematic Analysis
We transcribed the interaction sessions using Zoom’s transcription service. Our final dataset consisted of $5$ documents averaging $3049$ ($R=1642-4614$) words. We analyzed this data set with inductive thematic analysis [47]. We started with a single interview and went through the transcript one sentence at a time, assigning codes that captured the content of the sentence, or a set of related sentences. For each new interview we analyzed, we used existing codes generated based on previous interviews and assigned new codes only when there were no existing codes to capture the content of a specific statement or set of statements. Each time a new set of codes was added based on an interview, we revisited previously coded interviews to check if the new codes could be assigned to any statements in them. We found that we reached data saturation at $5$ interviews with only $2$ new codes being added in analyzing the final interview. We identified a set of $91$ content-based codes across all transcribed participant-radiologist interactions, and we grouped these codes into a set of $10$ broad themes that best captured the relationship between the codes.